OpenStudy (anonymous):

express 3/4 as an equivalent fraction having a denominator of 20

OpenStudy (anonymous):

20/4=5 3/4 = 3·5/4·5 = 15/20

OpenStudy (smurfy14):

15/20 what times 4=20? the answer is 5 so multiply 3 by 5 and you get 15/20